A thin copper rod 1.00 m long has a mass of 50.0 g. What is the minimum current in the rod that would allow it to levitate above the ground in a magnetic field of magnitude 0.100 T?
a.1.20 A
b.2.40 A
c.4.90 A
d.9.80 A
e.none of these answers
Answer #1
Concepts and reason

The concepts used to solve this problem are magnetic force and gravitational force.

First, equate the magnitude of the magnetic force and the gravitational force acting on the copper rod to determine the condition for levitation.

Use the relationship between the magnetic field acting on the rod, length of the rod, and the acceleration due to gravity to determine the required current in the rod.

Fundamentals

The expression for the magnetic force that acts on the copper rod is as follows:

Fm=BIL{F_m} = BIL

Here, the magnetic force is Fm{F_m}, the magnetic field is BB, the current in the copper wire is II, and the length of the wire is LL.

The expression for the gravitational force is as follows:

Fg=mg{F_g} = mg

Here, the gravitational force is Fg{F_g}, the mass of the copper rod is mm, the acceleration due to gravity is gg.

The expression for the condition for levitation of copper rod is as follows:

Fm=Fg\left| {{F_m}} \right| = \left| {{F_g}} \right|

Here, the magnitude of the magnetic force is Fm\left| {{F_m}} \right| and the magnitude of the gravitational force is Fg\left| {{F_g}} \right|

Levitation occurs when the magnetic force on the copper rod is balanced by the gravitational force acting on it.

The expression for this condition is as follows:

Fm=Fg\left| {{F_m}} \right| = \left| {{F_g}} \right|

Substitute BILBIL for Fm\left| {{F_m}} \right| and mgmg for Fg\left| {{F_g}} \right|.

BIL=mgBIL = mg

Rearrange the above expression for the current.

I=mgBLI = \frac{{mg}}{{BL}}

The expression for the current in the copper wire is as follows:

I=mgBLI = \frac{{mg}}{{BL}}

Substitute 50.0g50.0\,{\rm{g}} for mm, 9.8ms29.8\,{\rm{m}}{{\rm{s}}^{ - 2}} for gg, 0.100T0.100\,{\rm{T}} for BB, and 1.00m1.00\,{\rm{m}} for LL.

I=(50.0g)(1kg103g)(9.8ms2)(0.100T)(1.00m)=(50.0×103kg)(9.8ms2)(0.100T)(1.00m)=4.90A\begin{array}{c}\\I = \frac{{\left( {50.0\,{\rm{g}}} \right)\left( {\frac{{1\,{\rm{kg}}}}{{{{10}^3}\,{\rm{g}}}}} \right)\left( {9.8\,{\rm{m}}{{\rm{s}}^2}} \right)}}{{\left( {0.100\,{\rm{T}}} \right)\left( {1.00\,{\rm{m}}} \right)}}\\\\ = \frac{{\left( {50.0 \times {{10}^{ - 3}}\,{\rm{kg}}} \right)\left( {9.8\,{\rm{m}}{{\rm{s}}^2}} \right)}}{{\left( {0.100\,{\rm{T}}} \right)\left( {1.00\,{\rm{m}}} \right)}}\\\\ = 4.90\,{\rm{A}}\\\end{array}

The required current in the copper wire is I=4.90AI = 4.90\,{\rm{A}}.

Ans:

Thus, the correct option for the required current in the copper wire for it to levitate in the magnetic field is 4.90A4.90\,A.